OpenStudy (anonymous):

Determine whether the following series converges or diverges. If it converges, calculate the sum lower limit = 1 upper limit = infinity ((-3)^(n-1))(4^(-n))

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (-3)^{n-1}4^{-n}\]

OpenStudy (zarkon):

it is a geometric sum it converges

OpenStudy (zarkon):

\[\sum_{n=1}^{\infty} (-3)^{n-1}4^{-n}=\sum_{n=1}^{\infty} \frac{(-3)^{n-1}}{4^{n}}\] \[=\frac{1}{4}\sum_{n=1}^{\infty} \frac{(-3)^{n-1}}{4^{n-1}}\] \[=\frac{1}{4}\sum_{n=1}^{\infty} \left(-\frac{3}{4}\right)^{n-1}\]